Solve the equation.
9x = 3x^3
3x^3-9x=0{subtracted 9x from both sides}
3x(x^2-3){factored 3x out}
sqrt(x^2)=+-sqrt(3) {to take away ^2}
x=+-sqrt(3)

Found 2 solutions by algebrahouse.com, adrian438:
Answer by algebrahouse.com(1079)

Solve the equation.
9x = 3x^3
3x^3 - 9x = 0 {subtracted 9x from both sides}
3x(x^2 - 9) = 0 {factored 3x out}
3x(x + 3)(x - 3) = 0 {factored x2 - 9 into two binomials}
3x = 0 or x + 3 = 0 or x - 3 = 0 {set each factor equal to 0}
x = 0 or x = -3 or x = 3 {solved each equation for x}
x = 0,-3, or 3
www.algebrahouse.com

Answer by adrian438(1)

3x^3-9x=0{subtracted 9x from both sides}
3x(x^2-3){factored 3x out}
sqrt(x^2)=+-sqrt(3) {to take away ^2}
x=+-sqrt(3)
3x=0
x=0/3
x=0
answer: x=0,+-sqrt(3)